# Chicken Swarm Kernel Particle Filter: A Structured Rejuvenation Approach with KLD-Efficient Sampling

### v1

@@ 2.2 KLD-Based Adaptive Particle Sizing
Let the posterior approximation be represented by a histogram with $k$ non-empty bins. Fox’s KLD-sampling method Fox (2003) determines the number of samples $N$ required so that the Kullback–Leibler divergence between the empirical histogram and the (unknown) true posterior does not exceed a user-specified bound $\varepsilon$ with probability at least $1-\delta$. The resulting bound is
@@
if the particle cloud is narrow and concentrated, then $k$ is small and a relatively small $N$ suffices; if the particle cloud is broad and diverse, then $k$ is large and more particles are needed to satisfy the KLD constraint.
@@ 4.1 Setup and Notation
represent the corresponding bin probabilities. Given $N$ i.i.d. particles $X^{(1)},\dots,X^{(N)}\sim\mu$, denote by $N_{j}$ the number of particles in bin $B_{j}$, and define the number of occupied bins as:
@@ 4.3 From Contraction to Fewer Occupied Bins
For any distribution $\mu$ and sample size $N$, the expected number of occupied bins can be written explicitly as:
@@
Let $p,q\in[0,1]^{m}$ be two probability vectors with $p\succ q$. For any integer $N\geq 2$, let $K_{p}$ and $K_{q}$ be the numbers of occupied bins obtained from $N$ i.i.d. samples drawn from $p$ and $q$, respectively. Then:
@@ Appendix B Proof of Majorization Lemma
Given $N$ samples, the probability of a bin $j$ being empty is $(1-p_{j})^{N}$. Therefore, the probability of it being non-empty is $1-(1-p_{j})^{N}$.

### v2

# On the Interaction Between Chicken Swarm Rejuvenation and KLD-Adaptive Sampling in Particle Filters

@@ 2.2 KLD-Based Adaptive Particle Sizing
Let the posterior approximation be represented by a histogram with $k$ non-empty bins. Fox’s KLD-sampling method Fox (2003) determines the number of samples NN required so that the Kullback–Leibler divergence between the empirical histogram and the (unknown) true posterior does not exceed a user-specified bound $\varepsilon$ with probability at least $1-\delta$. The resulting bound is
@@
if the particle cloud is narrow and concentrated, then $k$ is small and a relatively small NN suffices; if the particle cloud is broad and diverse, then $k$ is large and more particles are needed to satisfy the KLD constraint.
@@ 4.1 Setup and Notation
represent the corresponding bin probabilities. Given NN i.i.d. particles $X^{(1)},\dots,X^{(N)}\sim\mu$, denote by $N_{j}$ the number of particles in bin $B_{j}$, and define the number of occupied bins as:
@@ 4.3 From Contraction to Fewer Occupied Bins
For any distribution $\mu$ and sample size NN, the expected number of occupied bins can be written explicitly as:
@@
Let $p,q\in[0,1]^{m}$ be two probability vectors with $p\succ q$. For any integer $N\geq 2$, let $K_{p}$ and $K_{q}$ be the numbers of occupied bins obtained from NN i.i.d. samples drawn from $p$ and $q$, respectively. Then:
@@ Appendix B Proof of Majorization Lemma
Given NN samples, the probability of a bin $j$ being empty is $(1-p_{j})^{N}$. Therefore, the probability of it being non-empty is $1-(1-p_{j})^{N}$.